OpenStudy (anonymous):

3. What is the excluded value of the rational expression 2x+6/4x-8 ? (1 point) a. –3 b.–2 c.0 d.2

OpenStudy (anonymous):

division by 0 is undefined, so solve for 0 in the numerator 4x-8=0 x=?

OpenStudy (anonymous):

i meant to say solve for 0 in the denominator***

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

is 4x-8 in the denominator or just 4x?

OpenStudy (anonymous):

\[\frac{2x+6}{4x-8}\] is the question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you want to find the value of x that makes the expression undefined, which is when the denominator equals 0. so solve for x: 4x-8=0

OpenStudy (anonymous):

It's D.